A high-jumper approaches the bar at 6.71m/s. exactly 83% of his kinetic energy is converted into potential energy due to his tremendous high-jumping form. Will he be able to clear the bar when it is set at 6'3"
1
Expert's answer
2019-04-08T09:24:03-0400

According to energy conservation law h = 0.83v2/2g = 0.83*6.712/19.6 = 1.907 (m), what is a little bit more than 6'3" = 1,905 (m). So, it will…